15) A 1.8 L volume of neon gas (monatomic) is at a pressure of 4.0 atmospheres and a temperature of 440 K. The atomic mass ofA 1.8 L volume of neon gas (monatomic) is at a pressure of 4.0 atmospheres and a temperature of 440 K. The atomic mass of neon is 20.2 g/mol. In this situation, the temperature of the gas is increased to 540 and the volume is increased to 3.8 L. The final pressure of the gas, in atmospheres, is closest to:

A) 2.1

B) 2.6

C) 3.0

D) 2.8

E) 2.3
Answer #1

The problem can be solved from ideal gas equation

Pi*Vi/To=Pf*Vf/Tf

Pi=initial pressure, Vi=initial volume,Ti=initial temperature

Pf=final pressure,Vf=final volume, Tf=final temperature

4*1.8/440=Pf*3.8/540

Pf=2.3 atmosphere

This is the required answer

Option E is correct